A Ferries wheel with a radius 14.0𝑚 is turning about the horizontal axis through its center. The linear speed of a passenger on the rim is constant and equal to 7.00 𝑚/𝑠. What are the magnitude and direction of the passenger’s acceleration as she passes through a) the lowest point in her circular motion? b) The highest point in her circular motion? c) How much time does it take the Ferries Wheel to make one revolution?


1
Expert's answer
2021-09-25T13:09:38-0400

Given:

"r=14.0\:\rm m"

"v=7.00\:\rm m/s"


a) the passenger’s acceleration the lowest point

"a=\frac{v^2}{r}=\frac{7.00^2}{14.0}=3.50\:\rm m/s^2\quad \rm (upright)"

b) the passenger’s acceleration the highest point


"a=\frac{v^2}{r}=\frac{7.00^2}{14.0}=3.50\:\rm m/s^2\quad \rm (downward)"

c) the period of rotation

"T=\frac{2\pi r}{v}=\frac{2*3.14*14.0}{7.00}=12.6\:\rm s"